### v1

# Introducing a Novel Subclass of Harmonic Functions with Close-to-Convex Properties

Abstract.
In this paper, we introduce a new subclass of close-to-convex harmonic functions. We present a sufficient coefficient condition for a function to be a member of this class. Furthermore, we establish a distortion theorem. These results lay the groundwork for extending the findings to function classes involving higher-order derivatives.
Key words and phrases:
Close-to-convex functions, Harmonic functions, Starlike functions, Distortion
2020 Mathematics Subject Classification:
30C45, 30C50
1. Introduction
In the realm of harmonic functions, every function $\mathfrak{f}$ belonging to the class $\mathcal{SH}^{0}$ can be expressed as $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}$, where
(1.1) $\mathfrak{u}(z)=z+\sum_{m=2}^{\infty}u_{m}z^{m},\quad\mathfrak{v}(z)=\sum_{m=2}^{\infty}v_{m}z^{m}$
with both functions being analytic in the open unit disk $\mathbb{E}=\left\{z\in\mathbb{C}:\left|z\right|<1\right\}$. Provided that $|\mathfrak{u}^{\prime}(z)|>|\mathfrak{v}^{\prime}(z)|$ in $\mathbb{E}$, $\mathfrak{f}$ is locally univalent and sense-preserving in $\mathbb{E}$. It’s noteworthy that when $\mathfrak{v}(z)$ is identically zero, $\mathcal{SH}^{0}$ contracts to class $\mathcal{S}$.
The subclasses of $\mathcal{S}$ that map $\mathbb{E}$ onto starlike and close-to-convex domains, respectively, are denoted by $\mathcal{S^{*}}$ and $\mathcal{K}$. Similarly, $\mathcal{SH}^{0,*}$ and $\mathcal{KH}^{0}$ represent subclasses of $\mathcal{SH}^{0}$ that map $\mathbb{E}$ onto their respective domains. (For further details, refer to [1, 3])
In 2005, Gao and Zhou [4] introduced the class
$\mathcal{K}_{s}=\left\{\mathfrak{\mathfrak{u}}\in\mathcal{S}:\mathrm{Re}\left\{\frac{z^{2}\mathfrak{u}^{\prime}(z)}{-\phi(z)\phi(-z)}\right\}>0\text{ for }z\in\mathbb{E}\right\}.$
In 2011, Şeker [5] introduced a class
$\mathcal{K}(k,\gamma)=\left\{\mathfrak{\mathfrak{u}}\in\mathcal{S}:\mathrm{Re}\left\{\frac{z^{k}\mathfrak{u}^{\prime}(z)}{\phi_{k}(z)}\right\}>\gamma,\ \ \ 0\leq\gamma<1\text{ and }z\in\mathbb{E}\right\}$
where $k\in\mathbb{Z}^{+},$ $\phi(z)=z+\sum_{m=2}^{\infty}c_{m}z^{m}\in S^{*}\left(\frac{k-1}{k}\right)$ and the definition of $\phi_{k}(z)$ is as follows:
(1.2) $\phi_{k}(z)=\prod_{v=0}^{k-1}\mu^{-v}\phi(\mu^{v}z)\quad(\mu^{k}=1)$
with $\mu=e^{2\pi i/k}$.
The classes $\mathcal{K}_{s}$ and $\mathcal{K}(k,\gamma)$, introduced by Gao and Zhou [4] and Şeker [5], respectively, focus solely on analytic functions and are independent of the variable $\overline{z}$. This restriction prevents the study of properties of harmonic functions that depend on the variable $\overline{z}$. This leads us to define a new function class that includes the co-analytic part of harmonic functions.
Observe that if $k=2$ in the class $\mathcal{K}(k,\gamma)$, the class $\mathcal{K}(\gamma)$ studied by Kowalczyk et al. [2] is obtained. For $k=2$ and $\gamma=0$, the class $\mathcal{K}_{s}$ studied by Gao and Zhou [4] is obtained. It is clear that the class $\mathcal{K}(k,\gamma)$ encompasses both classes. The class $\mathcal{KH}^{0}(k,\gamma)$, which we will define shortly, covers the class $\mathcal{K}(k,\gamma)$ since it is defined using harmonic conjugates of analytic functions belonging to the classes $\mathcal{K}(k,\gamma)$. Therefore, the class $\mathcal{KH}^{0}(k,\gamma)$ is a generalization of the $\mathcal{K}(k,\gamma)$, $\mathcal{K}_{s}$, and $\mathcal{K}(\gamma)$ classes.
This generalization allows for a broader investigation of geometric properties, such as distortion bounds and close-to-convexity, for harmonic functions that depend on $\overline{z}$.
Definition 1.1.
The class $\mathcal{KH}^{0}(k,\gamma)$ is defined as the collection of functions $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}\in\mathcal{SH}^{0}$ that adhere to the following inequality:
(1.3) $\mathrm{Re}\left\{\frac{z^{k}\mathfrak{u}^{\prime}(z)}{\phi_{k}(z)}-\gamma\right\}>\left|\frac{z^{k}\mathfrak{v}^{\prime}(z)}{\phi_{k}(z)}\right|,$
where $10\leq\gamma<$ 1 and $\phi_{k}(z)$ is given by (1.2).
Specifically, when $\mathfrak{v}(z)\equiv 0$, the class $\mathcal{KH}^{0}(k,\gamma)$ reduces to the class $\mathcal{K}(k,\gamma)$. Also, by setting $\mathfrak{v}(z)\equiv 0$, $k=2$ and $\gamma=0$, we obtain $\mathcal{KH}^{0}(2,0)=\mathcal{K}_{s}$. The inclusion of the co-analytic term $\overline{\mathfrak{v}(z)}$ extends these classes, providing a more general framework that accommodates both analytic and co-analytic components. Additionally, when $\phi(z)=z$, and by appropriately selecting the parameters, the class $\mathcal{KH}^{0}(k,\gamma)$ can be reduced to several well-known subclasses of harmonic functions, as outlined below.
  • i
    $\mathcal{KH}^{0}(k,0)=\mathcal{PH}^{0}$ [6].
  • ii
    $\mathcal{KH}^{0}(k,\gamma)=\mathcal{PH}^{0}(\gamma)$ [7, 8].
  • ii
    $\mathcal{KH}^{0}(k,0)=\mathcal{WH}^{0}(0)$ ([9]).
  • iv
    $\mathcal{KH}^{0}(k,\gamma)=\mathcal{WH}^{0}(0,\gamma)$ ([10]).
  • v
    $\mathcal{KH}^{0}(k,\gamma)=\mathcal{AH}^{0}(1,0,\gamma)$ ([11]).
  • vi
    $\mathcal{KH}^{0}(k,0)=\mathcal{RH}^{0}(0,0)$ ([12]).
For further details on harmonic function classes defined by differential inequality, see [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
In this work, we investigate the distortion theorems and coefficient bounds for functions in the class $\mathcal{KH}^{0}(k,\gamma)$ and demonstrate that functions within this class exhibit close-to-convex behavior.
2. Examples of Functions in the Class $\mathcal{KH}^{0}(k,\gamma)$
Example 2.1.
Let $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}=z+\frac{1-\gamma}{m}\overline{z}^{m}$ and $\phi(z)=z$. For $10\leq\gamma<$ 1 and $\left|z\right|<1$, we have
$\mathrm{Re}\left\{\frac{z^{k}\mathfrak{u}^{\prime}(z)}{\phi_{k}(z)}-\gamma\right\}$ $=1-\gamma>(1-\gamma)\left|z\right|^{m-1}=\left|\frac{z^{k}\mathfrak{v}^{\prime}(z)}{\phi_{k}(z)}\right|.$
Hence, $\mathfrak{f}\in\mathcal{KH}^{0}(k,\gamma)$.
The following examples can be given for the special case of the parameters in Example 2.1.
Example 2.2.
Let $\mathfrak{f}=z+\frac{99}{200}\overline{z}^{2}$, $\gamma=\frac{1}{100}$ and $\phi(z)=z$. Then $\mathfrak{f}\in\mathcal{KH}^{0}(k,\frac{33}{100}).$ The unit disk is mapped to a starlike region by the function $f$. The depiction in Figure 1 showcases the image of the set $\mathbb{E}$ under the transformation defined by $\mathfrak{f}(z)=z+\frac{99}{200}\overline{z}^{2}$.
Figure 1. Under the map $\mathfrak{f}=z+\frac{99}{200}\overline{z}^{2}$, the image of the unit disk.
Example 2.3.
Let $\mathfrak{f}=z+\frac{1}{10}\overline{z}^{2}$, $\gamma=\frac{4}{5}$, and $\phi(z)=z$. Then $\mathfrak{f}\in\mathcal{KH}^{0}(k,\frac{4}{5})$. The unit disk is mapped to a convex region by the function $f$. The depiction in Figure 2 showcases the image of the set $\mathbb{E}$ under the transformation defined by $\mathfrak{f}=z+\frac{1}{10}\overline{z}^{2}$.
Figure 2. Under the map $\mathfrak{f}=z+\frac{1}{10}\overline{z}^{2}$, the image of the unit disk.
Example 2.4.
Let $\mathfrak{f}=z+\frac{33}{100}\overline{z}^{3}$, $\gamma=\frac{1}{100}$ and $\phi(z)=z$. Then $\mathfrak{f}\in\mathcal{KH}^{0}(k,\frac{33}{100}).$ The unit disk is mapped to a starlike region by the function $f$. The depiction in Figure 3 showcases the image of the set $\mathbb{E}$ under the transformation defined by $\mathfrak{f}=z+\frac{33}{100}\overline{z}^{3}$.
Figure 3. Under the map $\mathfrak{f}=z+\frac{33}{100}\overline{z}^{3}$, the image of the unit disk.
Example 2.5.
Let $\mathfrak{f}=z+\frac{1}{15}\overline{z}^{3}$, $\gamma=\frac{4}{5}$ and $\phi(z)=z$. Then $\mathfrak{f}\in\mathcal{KH}^{0}(k,\frac{4}{5}).$ The unit disk is mapped to a convex region by the function $f$. The depiction in Figure 4 showcases the image of the set $\mathbb{E}$ under the transformation defined by $\mathfrak{f}=z+\frac{1}{15}\overline{z}^{3}$.
Figure 4. Under the map $\mathfrak{f}=z+\frac{1}{15}\overline{z}^{3}$, the image of the unit disk.
Example 2.6.
Let $\mathfrak{f}=z+\frac{99}{500}\overline{z}^{5}$, $\gamma=\frac{1}{100}$ and $\phi(z)=z$. Then $\mathfrak{f}\in\mathcal{KH}^{0}(k,\frac{99}{500}).$ The unit disk is mapped to a starlike region by the function $f$. The depiction in Figure 5 showcases the image of the set $\mathbb{E}$ under the transformation defined by $\mathfrak{f}=z+\frac{99}{500}\overline{z}^{5}$.
Figure 5. Under the map $\mathfrak{f}=z+\frac{99}{500}\overline{z}^{5}$, the image of the unit disk.
Example 2.7.
Let $\mathfrak{f}=z+\frac{1}{25}\overline{z}^{5}$, $\gamma=\frac{4}{5}$ and $\phi(z)=z$. Then $\mathfrak{f}\in\mathcal{KH}^{0}(k,\frac{4}{5}).$ The unit disk is mapped to a convex region by the function $f$. The depiction in Figure 6 showcases the image of the set $\mathbb{E}$ under the transformation defined by $\mathfrak{f}=z+\frac{1}{25}\overline{z}^{5}$.
Figure 6. Under the map $\mathfrak{f}=z+\frac{1}{25}\overline{z}^{5}$, the image of the unit disk.
3. Geometric Properties of the class $\mathcal{KH}^{0}(k,\gamma)$
First, we give a result that establishes a sufficient condition for $\mathfrak{f}\in\mathcal{SH}^{0}$ to be close-to-convex, which comes from Clunie and Sheil-Small [1].
Lemma 3.1.
Let $\mathfrak{u}$ and $\mathfrak{v}$ be analytic functions in $\mathbb{E}$, such that $|\mathfrak{v}^{\prime}(0)|<|\mathfrak{u}^{\prime}(0)|$, and for each $\varepsilon$ $(|\varepsilon|=1)$, $F_{\varepsilon}=\mathfrak{u}+\varepsilon\mathfrak{v}$ is close-to-convex. Then, $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}$ is close-to-convex in $\mathbb{E}$.
The result we will present now establishes a connection between the $\mathcal{KH}^{0}(k,\gamma)$ harmonic function class and the $\mathcal{K}(k,\gamma)$ analytic function class.
Theorem 3.2.
$\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}\in\mathcal{KH}^{0}(k,\gamma)$ if and only if $F_{\varepsilon}=\mathfrak{u}+\varepsilon\mathfrak{v}\in\mathcal{K}(k,\gamma)$ for each $\varepsilon\ (\left|\varepsilon\right|=1)$.
Proof.
Assume $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}\in\mathcal{KH}^{0}(k,\gamma)$. For each $\varepsilon\ (\left|\varepsilon\right|=1)$, we have
${\text{Re}}\left\{\frac{z^{k}{F_{\varepsilon}}^{\prime}(z)}{\phi_{k}(z)}\right\}$
$={\text{Re}}\left\{\frac{z^{k}{\mathfrak{u}}^{\prime}(z)}{\phi_{k}(z)}\right\}+\varepsilon{\text{Re}}\left\{\frac{z^{k}{\mathfrak{v}}^{\prime}(z)}{\phi_{k}(z)}\right\}$
$>{\text{Re}}\left\{\frac{z^{k}{\mathfrak{u}}^{\prime}(z)}{\phi_{k}(z)}\right\}-\left|\frac{z^{k}{\mathfrak{v}}^{\prime}(z)}{\phi_{k}(z)}\right|>\gamma\text{ }\left(z\in\mathbb{E}\right).$
Hence, $F_{\varepsilon}\in\mathcal{K}(k,\gamma)$ for each $\varepsilon\ (\left|\varepsilon\right|=1)$.
Conversely, suppose $F_{\varepsilon}=\mathfrak{u}+\varepsilon\mathfrak{v}\in\mathcal{K}(k,\gamma)$. Then,
${\text{Re}}\left\{\frac{z^{k}{\mathfrak{u}}^{\prime}(z)}{\phi_{k}(z)}\right\}>{\text{Re}}\left[-\varepsilon\frac{z^{k}{\mathfrak{v}}^{\prime}(z)}{\phi_{k}(z)}\right]+\gamma\text{ }\left(z\in\mathbb{E}\right).$
Choosing an appropriate $\varepsilon\ (\left|\varepsilon\right|=1)$ yields
${\text{Re}}\left\{\frac{z^{k}{\mathfrak{u}}^{\prime}(z)}{\phi_{k}(z)}-\gamma\right\}>\left|\frac{z^{k}{\mathfrak{v}}^{\prime}(z)}{\phi_{k}(z)}\right|\text{ }\left(z\in\mathbb{E}\right),$
and thus $\mathfrak{f}\in\mathcal{KH}^{0}(k,\gamma)$. ∎
We will now demonstrate that harmonic functions belonging to the class $\mathcal{KH}^{0}(k,\gamma)$ map the open unit disk onto a close-to-convex region. To this end, we first state the following lemma and subsequently establish that functions in the class $\mathcal{K}(k,\gamma)$ are close-to-convex within the open unit disk.
Lemma 3.3.
[4] Let $\phi(z)=z+\sum_{m=2}^{\infty}c_{m}z^{m}\in\mathcal{S^{*}}\left(\frac{k-1}{k}\right)$ for $k\geq 1$. Then,
(3.1) $\Phi_{k}(z)=\frac{\phi_{k}(z)}{z^{k-1}}=z+\sum_{m=2}^{\infty}C_{m}z^{m}\in S^{*}$
where $\phi_{k}(z)$ is given by (1.2).
Theorem 3.4.
If $F$ is a function in the class $\mathcal{K}(k,\gamma)$, then $F$ is close-to-convex of order $\gamma$ in the region $\mathbb{E}$.
Proof.
Let $F\in\mathcal{K}(k,\gamma)$. We have
${\text{Re}}\left\{\frac{z^{k}{F}^{\prime}(z)}{\phi_{k}(z)}\right\}={\text{Re}}\left\{\frac{z{F}^{\prime}(z)}{\Phi_{k}(z)}\right\}>\gamma,$
where $\Phi_{k}(z)$ is given by (3.1). Therefore, the function $F$ is close-to-convex of order $\gamma$ since $\Phi_{k}(z)\in\mathcal{S^{*}}$. ∎
Theorem 3.5.
Every function in the class $\mathcal{KH}^{0}(k,\gamma)$ is close-to-convex within the region $\mathbb{E}$.
Proof.
Let $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}$ belong to class $\mathcal{KH}^{0}(k,\gamma)$. Then by Theorem 3.2 the function $F_{\varepsilon}=\mathfrak{u}+\varepsilon\mathfrak{v}$ belongs to class $\mathcal{K}(k,\gamma)$ and by Theorem 3.4 also close to convex in $\mathbb{E}$. Therefore, by Lemma 3.1, $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}\in\mathcal{KH}^{0}(k,\gamma)$ is also close to convex in $\mathbb{E}$. ∎
In the following result, we derive a coefficient bound for functions belonging to the class $\mathcal{KH}^{0}(k,\gamma)$.
Theorem 3.6.
Let $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}\in\mathcal{KH}^{0}(k,\gamma)$. For $m\geq 2$, the following inequalities hold:
$\ \left|u_{m}\right|+\left|v_{m}\right|\leq\gamma+m(1-\gamma).$
For the function $\mathfrak{f}(z)=z+[\gamma+m(1-\gamma)]z^{m}$, every outcome is sharp and every equality is holds.
Proof.
Suppose that $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}\in\mathcal{KH}^{0}(k,\gamma)$. $F_{\varepsilon}=\mathfrak{u}+\varepsilon\mathfrak{v}\in\mathcal{K}(k,\gamma)$for $\varepsilon$ $\left(\left|\varepsilon\right|=1\right)$, according to Theorem 3.2. With respect to every $\varepsilon\ (\left|\varepsilon\right|=1),$ we possess
${\text{Re}}\left\{\frac{z^{k}{F_{\varepsilon}}^{\prime}(z)}{\phi_{k}(z)}\right\}={\text{Re}}\left\{\frac{z{F_{\varepsilon}}^{\prime}(z)}{\Phi_{k}(z)}\right\}={\text{Re}}\left\{\frac{z\left({\mathfrak{u}}^{\prime}(z)+\varepsilon\mathfrak{v}^{\prime}(z)\right)}{\Phi_{k}(z)}\right\}>\gamma.$
for $z\in\mathbb{E}$. On the other hand, there is an analytic function $\mathfrak{P}(z)=1+\underset{m=1}{\overset{\infty}{\sum}}\mathfrak{p}_{m}z^{m}$ in $\mathbb{E}$ whose real part is positive, satisfying
(3.2) $\frac{z\left({\mathfrak{u}}^{\prime}(z)+\varepsilon\mathfrak{v}^{\prime}(z)\right)}{\Phi_{k}(z)}=\gamma+\left(1-\gamma\right)\mathfrak{P}(z).$
or
(3.3) ${z\left({\mathfrak{u}}^{\prime}(z)+\varepsilon\mathfrak{v}^{\prime}(z)\right)}=\left[\gamma+\left(1-\gamma\right)\mathfrak{P}(z)\right]\Phi_{k}(z).$
Upon comparing the coefficients in (3.3), it can be observed that
(3.4) $m(u_{m}+\varepsilon v_{m})=C_{m}+(1-\gamma)\mathfrak{p}_{m-1}+(1-\gamma)\mathfrak{p}_{1}C_{m-1}+\dots+(1-\gamma)\mathfrak{p}_{m-2}C_{2}.$
Since $\Phi_{k}(z)$ is starlike, we have $\left|C_{m}\right|\leq m$, and since $Re\{\mathfrak{P}(z)\}>0$, we have $\left|\mathfrak{p}_{m}\right|\leq 2$ for $m\geq 1$. Hence, by equation 3.4, we have
(3.5) $m\left|u_{m}+\varepsilon v_{m}\right|\leq m\left[\gamma+m(1-\gamma)\right].$
Since $\varepsilon\left(\left|\varepsilon\right|=1\right)$ is arbitrary, it follows that the proof is concluded. The function $\mathfrak{f}(z)=z+$ $[\gamma+m(1-\gamma)]z^{m}$, demonstrates the sharpness of inequality. ∎
Now, we provide the necessary coefficient condition for a harmonic function to belong to the class $\mathcal{KH}^{0}(k,\gamma)$.
Theorem 3.7.
Let $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}\in\mathcal{SH}^{0}$ with the series expansions given by (1.1). If the following inequality holds:
(3.6) $\sum\limits_{m=2}^{\infty}2m\left(\left|u_{m}\right|+\left|v_{m}\right|\right)+\sum\limits_{m=2}^{\infty}\left(\left|1-2\gamma\right|+1\right)\left|C_{m}\right|\leq 2(1-\gamma),$
then $\mathfrak{f}\in\mathcal{KH}^{0}(k,\gamma)$.
Proof.
Consider $\mathfrak{u}$ and $\mathfrak{v}$ as functions with series expansions given by (1.1). Let $F_{\varepsilon}=\mathfrak{u}+\varepsilon\mathfrak{v}$ for each $\varepsilon\ (\left|\varepsilon\right|=1)$. Define the functions $\phi_{k}(z)$ and $\Phi_{k}(z)$ as given by (1.2) and (3.1), respectively. Then we can express $A$ as
$A$ $=$ $\left|zF_{\varepsilon}^{\prime}-\frac{\phi_{k}(z)}{z^{k-1}}\right|-\left|zF_{\varepsilon}^{\prime}+\frac{(1-2\gamma)\phi_{k}(z)}{z^{k-1}}\right|$
$=$ $\left|z(\mathfrak{u}+\varepsilon\mathfrak{v})^{\prime}-\Phi_{k}(z)\right|-\left|z(\mathfrak{u}+\varepsilon\mathfrak{v})^{\prime}+(1-2\gamma)\Phi_{k}(z)\right|$
$=$ $\left|\sum_{m=2}^{\infty}m(u_{m}+\varepsilon v_{m})z^{m}-\sum_{m=2}^{\infty}C_{m}z^{m}\right|$
-- $\left|(2-2\gamma)z+\sum_{m=2}^{\infty}m(u_{m}+\varepsilon v_{m})z^{m}+(1-2\gamma)\sum_{m=2}^{\infty}C_{m}z^{m}\right|$
$\leq$ $\sum_{m=2}^{\infty}m\left|u_{m}+\varepsilon v_{m}\right|\left|z\right|^{m}+\sum_{m=2}^{\infty}\left|C_{m}\right|\left|z\right|^{m}$
-- $\left((2-2\gamma)\left|z\right|-\sum_{m=2}^{\infty}m\left|u_{m}+\varepsilon v_{m}\right|\left|z\right|^{m}-\left|1-2\gamma\right|\sum_{m=2}^{\infty}\left|C_{m}\right|\left|z\right|^{m}\right)$
$<$ $\left\{-2(1-\gamma)+\sum_{m=2}^{\infty}2m\left|u_{m}+\varepsilon v_{m}\right|+\left(\left|1-2\gamma\right|+1\right)\sum_{m=2}^{\infty}\left|C_{m}\right|\right\}\left|z\right|$
Since $\varepsilon\left(\left|\varepsilon\right|=1\right)$ is arbitrary, and from the inequality (3.6), we obtain that $A<0$. This implies that $F_{\varepsilon}=\mathfrak{u}+\varepsilon\mathfrak{v}$ belongs to the class $\mathcal{K}(k,\gamma)$, and consequently, according to Theorem 2, it shows that $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}$ belongs to the class $\mathcal{KH}^{0}(k,\gamma)$. ∎
The result we present now provides the distortion bounds for functions in the class $\mathcal{KH}^{0}(k,\gamma)$.
Theorem 3.8.
Assuming $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}\in\mathcal{KH}^{0}(k,\gamma)$, the following inequalities hold for all $z$:
$\left|z\right|+\sum\limits_{m=2}^{\infty}(-1)^{m-1}[m(1-\gamma)+\gamma]\left|z\right|^{m}\leq\left|\mathfrak{f}(z)\right|\leq\left|z\right|+\sum\limits_{m=2}^{\infty}[m(1-\gamma)+\gamma]\left|z\right|^{m}.$
These are sharp inequality for the function $\mathfrak{f}(z)=z+\sum\limits_{m=2}^{\infty}[m(1-\gamma)+\gamma]z^{m}$.
Proof.
Let $\mathfrak{f}=\mathfrak{u}+\overline{\mathfrak{v}}\in\mathcal{KH}^{0}(k,\gamma).$ Then using Theorem 3.2, $F_{\varepsilon}=\mathfrak{u}+\varepsilon\mathfrak{v}\in\mathcal{K}(k,\gamma)$for each $\varepsilon\ (\left|\varepsilon\right|=1).$ Additionally, from Theorem 3.6 in [5], we obtain
(3.7) $\frac{1-(1-2\gamma)\left|z\right|}{(1+\left|z\right|)^{3}}\leq\left|F_{\varepsilon}^{\prime}(z)\right|\leq\frac{1+(1-2\gamma)\left|z\right|}{(1-\left|z\right|)^{3}}$
Since
$\left|F_{\varepsilon}^{\prime}(z)\right|$ $=$ $\left|\mathfrak{u}^{\prime}(z)+\varepsilon\mathfrak{v}^{\prime}(z)\right|$
$\leq$ $1+\sum\limits_{m=2}^{\infty}m[m(1-\gamma)+\gamma]\left|z\right|^{m-1}$
and
$\left|F_{\varepsilon}^{\prime}(z)\right|$ $=$ $\left|\mathfrak{u}^{\prime}(z)+\varepsilon\mathfrak{v}^{\prime}(z)\right|$
$\geq$ $1+\sum\limits_{m=2}^{\infty}(-1)^{m-1}m[m(1-\gamma)+\gamma]\left|z\right|^{m-1},$
in particular, we get
$\left|\mathfrak{u}^{\prime}(z)\right|+\left|\mathfrak{v}^{\prime}(z)\right|\leq 1+\sum\limits_{m=2}^{\infty}m[m(1-\gamma)+\gamma]\left|z\right|^{m-1}$
and
$\left|\mathfrak{u}^{\prime}(z)\right|-\left|\mathfrak{v}^{\prime}(z)\right|\geq 1+\sum\limits_{m=2}^{\infty}(-1)^{m-1}m[m(1-\gamma)+\gamma]\left|z\right|^{m-1}.$
Assume $\Gamma$ is the radial segment extending from 00 to $z$, so
[…]
$=$ $\left|z\right|+\sum\limits_{m=2}^{\infty}[m(1-\gamma)+\gamma]\left|z\right|^{m},$
and
$\left|\mathfrak{f}(z)\right|$$\geq$ $\int\limits_{\Gamma}\left(\left|\mathfrak{u}^{\prime}(\mathfrak{z})\right|-\left|\mathfrak{v}^{\prime}(\mathfrak{z})\right|\right)\left|d\mathfrak{z}\right|$
[…]
$=$ $\left|z\right|+\sum\limits_{m=2}^{\infty}(-1)^{m-1}[m(1-\gamma)+\gamma]\left|z\right|^{m}.$
Theorem 3.9.
The class $\mathcal{KH}^{0}(k,\gamma)$ is closed under convex combinations.
Proof.
Suppose $\mathfrak{f}_{\alpha}=\mathfrak{u}_{\alpha}+\overline{\mathfrak{v}_{\alpha}}\in\mathcal{KH}^{0}(k,\gamma)$ for $\alpha=1,2,...,p$ and $\sum\limits_{\alpha=1}^{p}s_{\alpha}=1$ $(0\leq s_{\alpha}\leq 1).$ The convex combination of functions $\mathfrak{f}_{\alpha}$ $\left(\alpha=1,2,...,p\right)$ can be expressed as:
$\mathfrak{f}(z)=\sum\limits_{\alpha=1}^{p}s_{\alpha}\mathfrak{f}_{\alpha}(z)=\mathfrak{u}(z)+\overline{\mathfrak{v}(z)},$
where
$\mathfrak{u}(z)=\sum\limits_{\alpha=1}^{p}s_{\alpha}\mathfrak{u}_{\alpha}(z)\text{ \ and \ }\mathfrak{v}(z)=\sum\limits_{\alpha=1}^{p}s_{\alpha}\mathfrak{v}_{\alpha}(z).$
Both $\mathfrak{u}$ and $\mathfrak{v}$ are analytic within the open unit disk $\mathbb{E}$, satisfying initial conditions $\mathfrak{u}(0)=\mathfrak{v}(0)=\mathfrak{u}^{\prime}(0)-1=\mathfrak{v}^{\prime}(0)=0$ and
${\text{Re}}\left[\frac{z^{k}\mathfrak{u}^{\prime}(z)}{\phi_{k}(z)}-\gamma\right]$ $=$ ${\text{Re}}\left[\sum\limits_{\alpha=1}^{p}s_{\alpha}\left(\frac{z^{k}\mathfrak{u}_{\alpha}^{\prime}(z)}{\phi_{k}(z)}-\gamma\right)\right]>\sum\limits_{\alpha=1}^{p}s_{\alpha}\left|\frac{z^{k}\mathfrak{v}_{\alpha}^{\prime}(z)}{\phi_{k}(z)}\right|=\left|\frac{z^{k}\mathfrak{v}^{\prime}(z)}{\phi_{k}(z)}\right|$
showing that $\mathfrak{f}\in\mathcal{KH}^{0}(k,\gamma)$. ∎
4. Conclusions
In this paper, we introduced a new class of harmonic functions denoted by $\mathcal{KH}^{0}(k,\gamma)$. We established a relationship between $\mathcal{KH}^{0}(k,\gamma)$ and $\mathcal{K}(k,\gamma)$. We demonstrated that $\mathcal{KH}^{0}(k,\gamma)$ is close-to-convex. For functions in the $\mathcal{KH}^{0}(k,\gamma)$ class, we derived coefficient bounds and distortion theorems. Finally, we proved that $\mathcal{KH}^{0}(k,\gamma)$ is closed under convolution.